Circular No. 6575

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) J. M. Veal and L. E. Snyder, University of Illinois; M. C. H. Wright,

J. R. Forster, W. Hoffman, M. Pound, I. de Pater, T. Helfer, R. L. Plambeck, G. Engargiola, and T. Wong, University of California at Berkeley; L. M. Woodney and M. F. A’Hearn, University of Maryland; P. Palmer, University of Chicago; and Y.-J. Kuan, Institute of Astronomy and Astrophysics, Academia Sinica, Taiwan, report the detection of HCO^{+} from comet C/1995 O1 using seven antennae of the Berkeley-Illinois-Maryland-Association Array in autocorrelation mode: “The J=1–0 transition at rest frequency 89.189 GHz was detected in emission on Feb. 19, 22, and 28 UT with peak antenna temperatures (T_{a}^{∗}) of 80, 80, and 100 mK, respectively. The HCO^{+} spectral lines are consistently asymmetric with a redward wing. The full widths at half maximum (not including the redward wing) are equal to 3.0 ± 0.2, 4.5 ± 0.2, and 3.0 ± 0.2 km/s, with line centers (also not including the redward wing) at +1.6 ± 0.2, +1.8 ± 0.2, and +1.1 ± 0.2 km/s with respect to the comet nucleus, respectively. Relative to the main emission component, the redward wing is ∼ 1/3 the peak temperature and is displaced by ∼ +3.7 km/s from the line center. We believe this to be the first detection of HCO^{+} in a comet.”

T. Kawabata and K. Ayani, Bisei Astronomical Observatory (BAO), report observations with K. Kinoshita and M. Fujii at the BAO 1-m telescope: “High-resolution (instrumental FWHM = 0.06 nm) slit spectra of C/1995 O1 show emission lines of Na D_{1} and D_{2} on Feb. 26.9 UT; the observed velocities within 7′′ of the comet’s nucleus in the north-south slit were −32 and −29 km/s for D_{1} and D_{2}, respectively, while nearby absorption lines have an average velocity of −48 ± 3 km/s. It appears that the Na emission has an asymmetric spatial profile and is more extended toward the south (dust-tail side) than the north.”

F. Mannucci, CAISMI, CNR, Florence; and G.-P. Tozzi, Osservatorio Astrofisico di Arcetri, write: “Near-infrared J, H, and K observations with the TIRGO telescope between Feb. 3 and 10 show several shells with a projected velocity of 0′′.80/hr, confirming a rotation period of ∼ 12 hr (cf. IAUC 6560). Preliminary analysis shows that the nucleus’ spin-axis position is almost perpendicular to the line-of-sight with p.a. ∼ 40^{o}; the latitude of the active spot (∼ −65^{o}) is more uncertain. The jet velocity appears to be on the order of 0.35–0.45 km/s.”

1997 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green